# Mitigating Omitted Variable Bias in Empirical Software Engineering

### v3

@@ 1 Introduction
Consider a common task of any empirical discipline: estimating the (average) effect that changing a variable $X$ has on another, dependent variable $Y$. In experimental study design parlance, $X$ is the treatment [^1] The term treatment comes from the practice of randomized controlled clinical trials, where it strictly denotes a binary variable (treatment vs. control group) that is randomly assigned. In statistical parlance, however, the term is routinely applied more generally to indicate any variable whose effect on an outcome is being analyzed. Under this general meaning, a treatment variable may not be randomly assigned [34] (for example, in an observational study) or it may be continuous [32] instead of binary. In this paper, we use “treatment” with this relaxed, more general meaning—as a synonym of “exposure”. or exposure, and $Y$ is the outcome or response. Here is a concrete example in the domain of software engineering, which we will look into more closely in Section 3: estimating how using a different programming language (treatment $X$) affects the quality (outcome $Y$) of a program implemented in that language. If we have observational data about $X$ and $Y$, we may simply fit a statistical model—anything from a simple linear regression to fancier devices—on these data, and interpret the fitted model’s parameters connecting $X$ and $Y$ as an estimate of this effect. In this process, a well-known snag is omitted variable bias: if there exists another unmeasured variable $Z$ that affects both $X$ and $Y$, our estimate of the effect $X\to Y$ will spuriously include the combined effect of $X$ and $Z$ instead of the effect of $X$ alone. Continuing our empirical software engineering example, $Z$ may be a programmer’s intrinsic skills, which are all but sure to also affect the quality $Y$ of the written programs. In such a scenario, additional data about $X$ and $Y$ (i.e., sampling more datapoints) is not going to help; in fact, it may just entrench our reliance on the biased estimate by reducing its variance and giving the false impression of reliability or “significance”.
Omitted variable bias is a widespread risk of any statistical analysis of observational data, regardless of whether one employs frequentist [35, 28, 71], Bayesian [45, 59, 21], or other kinds of machine learning models [2, 54]. In fact, it is always possible that some relevant variable was not measured, because it was unknown, inaccessible, or impractical, time-consuming, or expensive to measure with reasonable accuracy. This paper’s main contribution is presenting several mitigation strategies to cope with omitted variable bias, and demonstrating them in scenarios and examples that are relevant for software engineering empirical data analysis.
@@
Omitted variable bias is a common and relevant problem—especially for empirical software data. However, it is by no means the only pitfall of analyzing empirical data: as we further discuss in Section 2, there are plenty of other challenges such as the included variable bias [22], “precisely inaccurate” analyses that hide biases [51], and unrepresentative population samples [9]. While each challenge requires different measures, they all likely involve trade-offs between costs and benefits—similarly to the present paper’s outlook on dealing with omitted variable bias.
As we discuss more broadly in Section 2.1, there has been a growing interest in adopting robust, modern statistical practices in empirical software engineering research—especially those based on causal analysis. This paper contributes to this line of research, focusing on the practical and widespread problem of omitted variable bias.
@@ 1.1 Contributions
    replication package:  https://figshare.com/s/fe607d8eb7c4cedbac75[23]
@@ 1.2 Scope and Limitations
  • We focus on recovering causal effects, as opposed to optimizing predictive performance or model succinctness in a purely statistical setting; in related work, we demonstrated scenarios where these goals lead to different results [22].
  • The methods we present in the paper are well-suited for cross-sectional studies, where data from a population is collected at a specific point in time. These studies are very common in empirical software engineering. Section 2.5 outlines a few approaches that extend confounding analysis to time-dependent data, which arise in longitudinal or panel studies.
  • Omitted variable bias is only one of many different threats to the validity of empirical studies. Different kinds of confounding may require different statistical approaches [50]; more generally, other sources of bias (selection bias, measurement bias, …) may jeopardize a study’s design validity at different levels [82].
  • The DAG-based methods we consider in this paper have been gaining traction, as they support rigorous, yet intuitive foundations for causal modeling, and they are sufficiently general to accommodate a wide range of applications. However, other causal analysis frameworks exist—such as difference-in-differences, instrumental variables, regression discontinuity designs, matching methods, and synthetic control—that may be more suitable in certain contexts. Without going into details, a key difference of these approaches to causality is that they usually address confounding through preprocessing, design, or data structure (for example, relying on timing or instruments) rather than by modeling and adjusting within a statistical model as in DAG-based approaches [69]. A more detailed presentation of these alternatives is outside this paper’s scope, and is available elsewhere [38, 4].
  • The DAG-based methods we demonstrate in the paper rely on several technical assumptions including: the absence of causal loops (acyclicity of the DAG), probabilistic relations between variables (as opposed to purely deterministic relations, which can only be modeled indirectly [7]), and enough evidence to inform a plausible causal structural model. Section 2 presents the underlying methods in detail, compares them to other, related approaches, and outlines their relevance for empirical software engineering research.
@@ 1.3 Organization
Section 2 first presents relevant related work on the topics of mitigating biases in statistical analysis and, more generally, best statistical practices for empirical software engineering; then, it introduces the key notations and concepts that the rest of the paper relies on. The rest of the paper demonstrates how to deal with omitted variable bias in two illustrative examples: Section 3 uses a comparatively simpler model to estimate the effect of programming languages on code quality (also mentioned in the introduction), which serves as a relatively uncomplicated scenario. Then, Section 4 uses a more complex model to analyze the relation between team size (how many developers work on a project) and effort (how long it takes to complete the project). Although the two illustrative examples differ in complexity, they are both based on realistic scenarios and models that we investigated in our previous work [21, 18]. Section 5 serves as a high-level summary of the whole article, presenting guidelines, in the form of a sequence of analysis steps that generate fundamental information to support the design of a study that mitigates omitted variable bias. Finally, Section 6 concludes the paper with a short summary of the main contributions.
@@ 2 Related Work and Background
This section starts (in Section 2.1) with a brief overview of the origins of causal analysis techniques for observational data, and how they have been adopted in empirical sciences (including software engineering). Then, Section 2.2 and Section 2.3 introduce the key concepts (causal relations, DAGs, confounders) of causal analysis that we will develop in the rest of the paper. Finally, Section 2.4 positions the paper’s contributions by relating them to other forms of confounding and causal inference bias.
@@ 2.1 Background
One of science’s ultimate goals is understanding the processes that underlie observed phenomena. This means discovering cause/effect relations between variables, as opposed to mere statistical associations. While the roots of causal analysis date back to Neyman’s potential outcomes framework [68], a comprehensive understanding of causality has emerged only later, in the late part of the 20th century. The key milestones in the development of a robust understanding of causality include: 1. Rubin [63]built upon Neyman’s pioneering work, introducing a framework for causality in nonrandomized (observational) studies. 2. Angrist et al. [3]introduced instrumental variables [^2] In a nutshell, an instrumental variable is a variable that acts like a natural experiment on the treatment. based on the Neyman-Rubin framework. 3. Working at about the same time as Rubin, Pearl started focusing on structural (graph) models [57], which culminated in his celebrated techniques for rigorously analyzing causal effects and confounding Pearl [58].
@@
Among Pearl’s work, directed acyclic graphs (DAGs) have become widely used to model the structural dependencies between observed (and unobserved) variables. As we will demonstrate already with the simple example of Section 2.2, DAGs are, first of all, a practical notation to specify causal relations. [^3] An alternative approach to modeling causal relations is structural equation modeling (SEM). Semantically, Pearl’s causal DAG framework and SEM with latent variables are closely related. Specifically, Pearl’s original work built on SEM, and modern presentations of the causal DAG framework explicitly connect it to SEM [8]. Methodologically, however, the two approaches differ in their focus: causal DAGs (especially as we use them in the paper) specialize in modeling structural causal relations among (observed or latent) variables; SEM, instead, tends to mix measurement aspects (e.g., the relation between latent constructs and indicators) and structural ones (e.g., the relation between different constructs). They also support techniques to estimate the strength of the causal relation among some nodes in the graph. Usually this is done by constructing a (linear) statistical model among variables, selected according to the DAG’s structure. (We will demonstrate this shortly in Section 2.3.) Nowadays, causal analysis based on DAGs is routinely used in disciplines with a strong empirical component such as medicine [81, 70], epidemiology [27], economics [37], and biology [43].
@@
A key question when working with DAGs is how to build a realistic DAG in the first place. In fields where the underlying fundamental mechanisms are well understood, a DAG can be built based on expert knowledge and previous work. Another approach is causal discovery (also called structural learning), which tries to identify causal relations from data [67]. While causal discovery algorithms have made significant progress in the last decade [84]—in part on the wave of the recent machine learning boom—they remain heuristic approaches that work correctly only under precise assumptions about the possible interactions. This limitation is intrinsic, as one of the fundamental results of Pearl’s framework is that causal relations cannot be inferred from data alone (at least not without limiting assumptions). Regardless of whether they come from expert knowledge, are inferred heuristically from data, or simply encode some (plausible) hypotheses, DAGs remain a practical tool to precisely denote, validate, and reason about the causal relations in a system.
@@
While causal analysis techniques are not widely used in software engineering empirical research, they are gradually gaining traction. Siebert [66]’s recent survey reports 31 studies in empirical software engineering that targeted some kind of causal analysis technique. All of the reviewed papers were published in the last 15 years, which confirms that causal analysis is not yet an established practice but is slowly gaining popularity. The majority (17 out of 31) of papers reviewed by Siebert are about fault localization and refer to Baah et al. [5]—the first contribution that tried to apply a causal view instead of the traditional, purely correlational analysis that is commonplace in fault localization techniques [83, 85, 61]. Testing is another area targeted by several of the studies reviewed by Siebert; these include applications to mutation testing [44], simulation testing [12], and A/B testing [47]. The other papers reviewed by Siebert target various topics such as performance analysis (in one case still linked to fault localization [65]).
@@
In the last few years, some more empirical software engineering research was published, targeting varied topics such as: 1. modeling rules of human knowledge and how they are made available to artificial intelligence systems [31]; 2. studying the impact of social media posts on the popularity of open-source projects [17]; 3. analyzing dependencies in configurable software systems [29]; 4. studying the impact of programming languages on coding competitions [22].
@@
More broadly, these techniques belong to a broad and growing body of work on modernizing statistical practices [50]. This work has been the response to the realization that some “traditional” and widely used statistical methods have limitations in terms of interpretability and real-world significance [79, 1, 53]. While improving statistical practices is only a piece of the puzzle, inadequate methods have certainly concurred to the issues with replicability experienced in empirical research areas as diverse as medicine [39], psychology [14], and economics [10]. Since empirical software engineering research shares challenges and practices with these domains, adopting more robust data analysis practices will benefit the quality and impact of our research field [25, 42, 48, 55, 41].
@@ 2.2 Causal Dependencies and DAGs
Process $p_{1}$ in 1(a) produces values of $X$ that are drawn randomly from a normal distribution with zero mean and unit standard deviation; and values of $Y$ that are linearly proportional to those of $X$. Even in such an ideal scenario, any empirical measure of $X$ and $Y$ will include some measurement error $\epsilon$, which process $p_{1}$ models as another normal random variable that, together with $X$, determines the value of $Y$. We stress that we interpret 1(a)’s equations as capturing the causal dependencies among variables: $X$ and $\epsilon$ are drawn randomly (and independent of each other), and their random values determine (“cause”) the value of $Y$ in each draw. Correspondingly, the DAG (directed acyclic graph [58]) in 1(a) captures this causal relation between $X$ and $Y$ in a qualitative way: an edge connects $X$ to $Y$ to denote that the values of $X$ and $Y$ are related; furthermore, the edge is directed from $X$ to $Y$ to denote that changing $X$ directly affects $Y$, that is, it causes $Y$ to change.
@@ 2.3 Inference and Confounders
As we will see in our examples, correcting for possible confounders is not always possible or practical. On the one hand, some confounding variables may simply be inaccessible; on the other hand, the model of causality that underlies the application of adjustment techniques may itself be imprecise or uncertain. In such cases, a sensitivity analysis can mitigate uncertainty and clarify the boundary of validity of our findings. In a nutshell, a sensitivity analysis is a sort of “what if” analysis that quantifies the robustness of the results under different assumptions about the underlying process. [^6] It is important to notice that, despite their superficial similarities, the practice of sensitivity analysis and the malpractice of data dredging (also known as “$p$-hacking”) have opposite goals: “In $p$-hacking, many justifiable analyses are tried, and the one that attains statistical significance is reported. In sensitivity analysis, many justifiable analyses are tried, and all of them are described.” [50, p. 319]. A sensitivity analysis can corroborate results and expose limits of validity, but is not expected to give any kind of definitive, binary answer.

### v4

# 22email: bugcounting.net

@@ 1 Introduction
Consider a common task of any empirical discipline: estimating the (average) effect that changing a variable $X$ has on another, dependent variable $Y$. In experimental study design parlance, $X$ is the treatment [^1] The term treatment comes from the practice of randomized controlled clinical trials, where it strictly denotes a binary variable (treatment vs. control group) that is randomly assigned. In statistical parlance, however, the term is routinely applied more generally to indicate any variable whose effect on an outcome is being analyzed. Under this general meaning, a treatment variable may not be randomly assigned Holland (1986) (for example, in an observational study) or it may be continuous Hirano and Imbens (2004) instead of binary. In this paper, we use “treatment” with this relaxed, more general meaning—as a synonym of “exposure”. or exposure, and $Y$ is the outcome or response. Here is a concrete example in the domain of software engineering, which we will look into more closely in section 3: estimating how using a different programming language (treatment $X$) affects the quality (outcome $Y$) of a program implemented in that language. If we have observational data about $X$ and $Y$, we may simply fit a statistical model—anything from a simple linear regression to fancier devices—on these data, and interpret the fitted model’s parameters connecting $X$ and $Y$ as an estimate of this effect. In this process, a well-known snag is omitted variable bias: if there exists another unmeasured variable $Z$ that affects both $X$ and $Y$, our estimate of the effect $X\to Y$ will spuriously include the combined effect of $X$ and $Z$ instead of the effect of $X$ alone. Continuing our empirical software engineering example, $Z$ may be a programmer’s intrinsic skills, which are all but sure to also affect the quality $Y$ of the written programs. In such a scenario, additional data about $X$ and $Y$ (i.e., sampling more datapoints) is not going to help; in fact, it may just entrench our reliance on the biased estimate by reducing its variance and giving the false impression of reliability or “significance”.
Omitted variable bias is a widespread risk of any statistical analysis of observational data, regardless of whether one employs frequentist (Holt et al., 2014; Gren et al., 2017; Torkar et al., 2022), Bayesian (Levén et al., 2024; Penzenstadler et al., 2022; Furia et al., 2022), or other kinds of machine learning models (Andaur Navarro et al., 2021; Mehrabi et al., 2021). In fact, it is always possible that some relevant variable was not measured, because it was unknown, inaccessible, or impractical, time-consuming, or expensive to measure with reasonable accuracy. This paper’s main contribution is presenting several mitigation strategies to cope with omitted variable bias, and demonstrating them in scenarios and examples that are relevant for software engineering empirical data analysis.
@@
Omitted variable bias is a common and relevant problem—especially for empirical software data. However, it is by no means the only pitfall of analyzing empirical data: as we further discuss in section 2, there are plenty of other challenges such as the included variable bias (Furia et al., 2023), “precisely inaccurate” analyses that hide biases (McFarland and McFarland, 2015), and unrepresentative population samples (Bradley et al., 2021). While each challenge requires different measures, they all likely involve trade-offs between costs and benefits—similarly to the present paper’s outlook on dealing with omitted variable bias.
As we discuss more broadly in subsection 2.1, there has been a growing interest in adopting robust, modern statistical practices in empirical software engineering research—especially those based on causal analysis. This paper contributes to this line of research, focusing on the practical and widespread problem of omitted variable bias.
@@ 1.1 Contributions
    replication package:  https://figshare.com/s/fe607d8eb7c4cedbac75Furia and Torkar (2025)
@@ 1.2 Scope and Limitations
  • We focus on recovering causal effects, as opposed to optimizing predictive performance or model succinctness in a purely statistical setting; in related work, we demonstrated scenarios where these goals lead to different results (Furia et al., 2023).
  • The methods we present in the paper are well-suited for cross-sectional studies, where data from a population is collected at a specific point in time. These studies are very common in empirical software engineering. For comparison, subsection 2.5 outlines a few approaches that extend confounding analysis to time-dependent data, which arise in longitudinal or panel studies.
  • Omitted variable bias is only one of many different threats to the validity of empirical studies. Different kinds of confounding may require different statistical approaches (McElreath, 2020); more generally, other sources of bias (selection bias, measurement bias, …) may jeopardize a study’s design validity at different levels Wohlin et al. (2012).
  • The DAG-based methods we consider in this paper have been gaining traction, as they support rigorous, yet intuitive foundations for causal modeling, and they are sufficiently general to accommodate a wide range of applications. However, other causal analysis frameworks exist—such as difference-in-differences, instrumental variables, regression discontinuity designs, matching methods, and synthetic control—that may be more suitable in certain contexts. Without going into details, a key difference of these approaches to causality is that they usually address confounding through preprocessing, design, or data structure (for example, relying on timing or instruments) rather than by modeling and adjusting within a statistical model as in DAG-based approaches (Stuart, 2010). A more detailed presentation of these alternatives is outside this paper’s scope, and is available elsewhere (Imbens and Wooldridge, 2009; Athey and Imbens, 2017).
  • The DAG-based methods we demonstrate in the paper rely on several technical assumptions including: the absence of causal loops (acyclicity of the DAG), probabilistic relations between variables (as opposed to purely deterministic relations, which can only be modeled indirectly Berrie et al. (2025)), and enough evidence to inform a plausible causal structural model. section 2 presents the underlying methods in detail, compares them to other, related approaches, and outlines their relevance for empirical software engineering research.
@@ 1.3 Organization
section 2 first presents relevant related work on the topics of mitigating biases in statistical analysis and, more generally, best statistical practices for empirical software engineering; then, it introduces the key notations and concepts that the rest of the paper relies on. The rest of the paper demonstrates how to deal with omitted variable bias in two illustrative examples: section 3 uses a comparatively simpler model to estimate the effect of programming languages on code quality (also mentioned in the introduction), which serves as a relatively uncomplicated scenario. Then, section 4 uses a more complex model to analyze the relation between team size (how many developers work on a project) and effort (how long it takes to complete the project). Although the two illustrative examples differ in complexity, they are both based on realistic scenarios and models that we investigated in our previous work (Furia et al., 2022; Feldt et al., 2025). section 5 serves as a high-level summary of the whole article, presenting guidelines, in the form of a sequence of analysis steps that generate fundamental information to support the design of a study that mitigates omitted variable bias. Finally, section 6 concludes the paper with a short summary of the main contributions.
@@ 2 Related Work and Background
This section starts (in subsection 2.1) with a brief overview of the origins of causal analysis techniques for observational data, and how they have been adopted in empirical sciences (including software engineering). Then, subsection 2.2 and subsection 2.3 introduce the key concepts (causal relations, DAGs, confounders) of causal analysis that we will develop in the rest of the paper. Finally, subsection 2.4 positions the paper’s contributions by relating them to other forms of confounding and causal inference bias.
@@ 2.1 Background
One of science’s ultimate goals is understanding the processes that underlie observed phenomena. This means discovering cause/effect relations between variables, as opposed to mere statistical associations. While the roots of causal analysis date back to Neyman’s potential outcomes framework (Splawa-Neyman et al., 1990), a comprehensive understanding of causality has emerged only later, in the late part of the 20th century. The key milestones in the development of a robust understanding of causality include: 1. Rubin (1974)built upon Neyman’s pioneering work, introducing a framework for causality in nonrandomized (observational) studies. 2. Angrist et al. (1996)introduced instrumental variables [^2] In a nutshell, an instrumental variable is a variable that acts like a natural experiment on the treatment. based on the Neyman-Rubin framework. 3. Working at about the same time as Rubin, Pearl started focusing on structural (graph) models (Pearl, 1982), which culminated in his celebrated techniques for rigorously analyzing causal effects and confounding Pearl (2009).
@@
Among Pearl’s work, directed acyclic graphs (DAGs) have become widely used to model the structural dependencies between observed (and unobserved) variables. As we will demonstrate already with the simple example of subsection 2.2, DAGs are, first of all, a practical notation to specify causal relations. [^3] An alternative approach to modeling causal relations is structural equation modeling (SEM). Semantically, Pearl’s causal DAG framework and SEM with latent variables are closely related. Specifically, Pearl’s original work built on SEM, and modern presentations of the causal DAG framework explicitly connect it to SEM (Bollen and Pearl, 2013). Methodologically, however, the two approaches differ in their focus: causal DAGs (especially as we use them in the paper) specialize in modeling structural causal relations among (observed or latent) variables; SEM, instead, tends to mix measurement aspects (e.g., the relation between latent constructs and indicators) and structural ones (e.g., the relation between different constructs). They also support techniques to estimate the strength of the causal relation among some nodes in the graph. Usually this is done by constructing a (linear) statistical model among variables, selected according to the DAG’s structure. (We will demonstrate this shortly in subsection 2.3.) Nowadays, causal analysis based on DAGs is routinely used in disciplines with a strong empirical component such as medicine (Williams et al., 2018; Tennant et al., 2021), epidemiology (Greenland et al., 1999), economics (Imbens, 2020), and biology (Laubach et al., 2021).
@@
A key question when working with DAGs is how to build a realistic DAG in the first place. In fields where the underlying fundamental mechanisms are well understood, a DAG can be built based on expert knowledge and previous work. Another approach is causal discovery (also called structural learning), which tries to identify causal relations from data (Spirtes and Zhang, 2016). While causal discovery algorithms have made significant progress in the last decade (Zanga et al., 2022)—in part on the wave of the recent machine learning boom—they remain heuristic approaches that work correctly only under precise assumptions about the possible interactions. This limitation is intrinsic, as one of the fundamental results of Pearl’s framework is that causal relations cannot be inferred from data alone (at least not without limiting assumptions). Regardless of whether they come from expert knowledge, are inferred heuristically from data, or simply encode some (plausible) hypotheses, DAGs remain a practical tool to precisely denote, validate, and reason about the causal relations in a system.
@@
While causal analysis techniques are not widely used in software engineering empirical research, they are gradually gaining traction. Siebert (2022)’s recent survey reports 31 studies in empirical software engineering that targeted some kind of causal analysis technique. All of the reviewed papers were published in the last 15 years, which confirms that causal analysis is not yet an established practice but is slowly gaining popularity. The majority (17 out of 31) of papers reviewed by Siebert are about fault localization and refer to Baah et al. (2010)—the first contribution that tried to apply a causal view instead of the traditional, purely correlational analysis that is commonplace in fault localization techniques Wong et al. (2016); Zou et al. (2021); Rezaalipour and Furia (2024). Testing is another area targeted by several of the studies reviewed by Siebert; these include applications to mutation testing Lee et al. (2021), simulation testing (Clark et al., 2022), and A/B testing (Liu et al., 2022). The other papers reviewed by Siebert target various topics such as performance analysis (in one case still linked to fault localization (Scholz and Torkar, 2021)).
@@
In the last few years, some more empirical software engineering research was published, targeting varied topics such as: 1. modeling rules of human knowledge and how they are made available to artificial intelligence systems (Heyn and Knauss, 2022); 2. studying the impact of social media posts on the popularity of open-source projects (Fang et al., 2022); 3. analyzing dependencies in configurable software systems (Halpern, 2015); 4. studying the impact of programming languages on coding competitions (Furia et al., 2023).
@@
More broadly, these techniques belong to a broad and growing body of work on modernizing statistical practices (McElreath, 2020). This work has been the response to the realization that some “traditional” and widely used statistical methods have limitations in terms of interpretability and real-world significance Wasserstein and Lazar (2016); Amrehin et al. (2019); McShane et al. (2019). While improving statistical practices is only a piece of the puzzle, inadequate methods have certainly concurred to the issues with replicability experienced in empirical research areas as diverse as medicine Ioannidis (2005), psychology Collaboration (2015), and economics Camerer et al. (2016). Since empirical software engineering research shares challenges and practices with these domains, adopting more robust data analysis practices will benefit the quality and impact of our research field González-Barahona and Robles (2012); Krishnamurthi and Vitek (2015); Madeyski and Kitchenham (2017); Menzies and Shepperd (2019); Jørgensen et al. (2016).
@@ 2.2 Causal Dependencies and DAGs
Process $p_{1}$ in 1(a) produces values of $X$ that are drawn randomly from a normal distribution with zero mean and unit standard deviation; and values of $Y$ that are linearly proportional to those of $X$. Even in such an ideal scenario, any empirical measure of $X$ and $Y$ will include some measurement error $\epsilon$, which process $p_{1}$ models as another normal random variable that, together with $X$, determines the value of $Y$. We stress that we interpret 1(a)’s equations as capturing the causal dependencies among variables: $X$ and $\epsilon$ are drawn randomly (and independent of each other), and their random values determine (“cause”) the value of $Y$ in each draw. Correspondingly, the DAG (directed acyclic graph (Pearl, 2009)) in 1(a) captures this causal relation between $X$ and $Y$ in a qualitative way: an edge connects $X$ to $Y$ to denote that the values of $X$ and $Y$ are related; furthermore, the edge is directed from $X$ to $Y$ to denote that changing $X$ directly affects $Y$, that is, it causes $Y$ to change.
@@ 2.3 Inference and Confounders
As we will see in our examples, correcting for possible confounders is not always possible or practical. On the one hand, some confounding variables may simply be inaccessible; on the other hand, the model of causality that underlies the application of adjustment techniques may itself be imprecise or uncertain. In such cases, a sensitivity analysis can mitigate uncertainty and clarify the boundary of validity of our findings. In a nutshell, a sensitivity analysis is a sort of “what if” analysis that quantifies the robustness of the results under different assumptions about the underlying process. [^6] It is important to notice that, despite their superficial similarities, the practice of sensitivity analysis and the malpractice of data dredging (also known as “$p$-hacking”) have opposite goals: “In $p$-hacking, many justifiable analyses are tried, and the one that attains statistical significance is reported. In sensitivity analysis, many justifiable analyses are tried, and all of them are described.” (McElreath, 2020, p. 319). A sensitivity analysis can corroborate results and expose limits of validity, but is not expected to give any kind of definitive, binary answer.